### v1

# Masked Autoencoder Self Pre-Training for Defect Detection in Microelectronics

Abstract
Whereas in general computer vision, transformer-based architectures have quickly become the gold standard, microelectronics defect detection still heavily relies on convolutional neural networks (CNNs). We hypothesize that this is due to the fact that a) transformers have an increased need for data and b) labelled image generation procedures for microelectronics are costly, and labelled data is therefore sparse. Whereas in other domains, pre-training on large natural image datasets can mitigate this problem, in microelectronics transfer learning is hindered due to the dissimilarity of domain data and natural images. Therefore, we evaluate self pre-training, where models are pre-trained on the target dataset, rather than another dataset. We propose a vision transformer (ViT) pre-training framework for defect detection in microelectronics based on masked autoencoders (MAE). In MAE, a large share of image patches is masked and reconstructed by the model during pre-training. We perform pre-training and defect detection using a dataset of less than 10.000 scanning acoustic microscopy (SAM) images labelled using transient thermal analysis (TTA). Our experimental results show that our approach leads to substantial performance gains compared to a) supervised ViT, b) ViT pre-trained on natural image datasets, and c) state-of-the-art CNN-based defect detection models used in the literature. Additionally, interpretability analysis reveals that our self pre-trained models, in comparison to ViT baselines, correctly focus on defect-relevant features such as cracks in the solder material. This demonstrates that our approach yields fault-specific feature representations, making our self pre-trained models viable for real-world defect detection in microelectronics.
Keywords:
Masked Autoencoder Vision Transformer Pretraining Microelectronics LED Solder Joint Quality Control
1 Introduction
Figure 1: Masked Autoencoder Framework. In MAE pre-training, a large share (e.g. $75\%$) of input patches is masked. The encoder (ENC) then performs self-attention for visible patches only. After adding mask-tokens, which serve as placeholders for missing inputs, the decoder (DEC) reconstructs the complete image. Thereby, the model learns meaningful representations of input data.
Reliable solder joints are crucial for the continued miniaturization and enhanced functionality of microelectronics, with applications spanning consumer electronics, automotive systems, healthcare, and defense [25]. Defect detection in microelectronics often relies on images, which are obtained by intricate procedures like scanning microscopy or X-ray imaging. Image data are then used in smart manufacturing processes such as Automated Optical Inspection (AOI) to ensure product reliability and performance [1, 31].
In particular, deep learning based vision models have recently shown strong performance in AOI tasks like defect detection [41, 42, 27, 18]. While convolutional neural networks (CNNs) have traditionally been the dominant choice for deep learning based computer vision, vision transformer (ViT) models have recently emerged as the gold standard in the field [20, 22, 35]. However, the application of transformer models to defect detection in microelectronics remains underexplored, with the field often relying on conventional CNNs and recent survey studies sometimes not mentioning transformer models at all [27, 18]. One main reason for this could be that training transformers from scratch usually requires a large amount of data [40, 13, 46]. In microelectronics manufacturing - especially for microscale solder joints - labelled image data collection often relies on expensive procedures, resulting in small and imbalanced datasets [1]. Additionally, fine-tuning transformers pre-trained on large datasets usually requires at least some similarity between the pre-training domain and the target domain [2]. This circumstance is problematic for microelectronics since most pre-trained transformers are trained on natural image datasets like ImageNet [11] that are highly dissimilar from target datasets in this domain (see Figure 2).
Inspired by other domains facing similar challenges, such as healthcare, this paper explores the potential of a pure transformer model for microelectronics defect detection by leveraging self pre-training, where models are pre-trained directly on the target dataset rather than an extensive natural image dataset [50]. Specifically, we use masked autoencoders (MAEs) [16], which mask a large share of image patches and task the model with reconstructing the missing inputs (see Figure 1). MAEs are resource-efficient and well-suited for smaller datasets, as they do not require large batch sizes like contrastive pre-training approaches. Additionally, the repeated randomized masking of different image patches presents the autoencoder with diverse reconstruction tasks, even with limited data. Using a dataset of less than 10.000 scanning acoustic microscopy (SAM) images of microscale LED solder joints labelled using transient thermal analysis (TTA), we compare our approach to a) purely supervised ViT, b) ViT pre-trained on a natural image dataset, and c) state-of-the-art CNN architectures proposed in the literature on industrial defect detection. In particular, we fine-tune our models for a regression task where models predict how far a given LED is from failure.
1.0.1 Contributions
Our experimental findings indicate that, with our approach, models learn generalizable representations in self pre-training that yield substantial performance increases in defect detection for microelectronics. In particular, our model outperforms a) purely supervised ViTs by $21.1\%$, b) ViTs pre-trained on Imagenet by $10.2\%$ and c) state-of-the-art CNN defect detection models by $5.3\%$. Also, we find that our models learn fault-specific representations, focusing on actual damages in the solder material when performing defect detection, while other ViT models have scattered attention or learn shortcuts. In addition, due to the simplicity and low data requirements of our framework, it can be easily integrated into real-world inline quality control procedures such as AOI.
(a) Natural image data from ImageNet [11].
(b) Scanning Acoustic Microscopy images from our target dataset [31].
Figure 2: Domain gap between natural images and images from our target domain.
2 Related Work
2.0.1 Vision Transformer
ViT [13] is a vision adaption of the self-attention based transformer architecture [40], which constitutes the backbone of large language models. ViT has demonstrated effectiveness across various vision tasks, including image classification, object detection, and segmentation [19, 33, 37]. Related research on ViT includes domain-specific fine-tuning of foundational models and architectural adaptions that aim at improved domain generalization or increased data efficiency: Alijiani et al. [2] discuss various ViT domain adaption and generalization approaches on the feature extraction level, the instance level and the model architecture level. Papa et al. [26] compare ViT approaches aiming at increased data efficiency, such as Data Efficient Transformers (DeiT) [38, 39] or Shifted Window Transformers (Swin) [21]. Since we leverage representations learned in pre-training, our approach is similar to domain specific fine-tuning. However, our framework differs from such approaches since we pre-train ViT from scratch on the target dataset, rather than using foundational models pre-trained on large natural image datasets.
2.0.2 Self-Supervised Learning
SSL aims to learn meaningful representations from data without labels. Instead of labels, intrinsic relationships of the data serve as guidance for the training process. Balestriero et al. [3] distinguish four sub-types of SSL approaches: (1) deep metric learning approaches such as SimCLR [10] or NNCLR [14], where models recognize similarity in differently transformed inputs, (2) self-distillation learning approaches like BYOL [15], SimSIAM [9] or DINO [7], where different input transforms serve as inputs into two separate encoders before letting one serve as a predictor of the other’s output, (3) canonical correlation analysis approaches such as VICReg [5], BarlowTwins [45] or SWAV [6], which are founded on the principle of understanding relationships between two variables by examining their cross-covariance matrices and (4) masked image modelling (MIM) approaches like SimMIM [43], where images undergo partial masking and models reconstruct the missing inputs. We follow the MIM approach due to its simplicity, efficacy, and low data requirements.
[…]
2.0.3 Microelectronics Quality Control
In microelectronics research and manufacturing, image-based quality control procedures like AOI are standard practice [1]. Deep learning based approaches have significantly advanced the field in recent years, including but not limited to the subdomain of solder joints [41, 42, 27, 18]. Samavatian et al. [29] present an iterative machine learning framework that enhances the accuracy of solder joint lifetime prediction by utilizing a self-healing dataset iteratively injected into a correlation-driven neural network (CDNN). Salameh et al. [28] demonstrate the use of deep neural networks, combined with finite element simulations, as a rapid and comprehensive tool for solder joint reliability analysis under mechanical loading. Muench et al. [24] propose a methodology for predicting damage progression in solder contacts and compare an MLP network with a long short-term memory (LSTM) model using production-like synthetic data. Zhang et al. [48] employ various deep learning models, including CNN and LSTM models, for automatic solder joint defect detection using X-ray images. Zhang et al. [49] combine CNN and transformer components in a model intended for printed control board solder joints. However, Zhang et al. use transformer blocks only as small parts of their network aided by convolutional layers and report poor performances for pure transformer models [49]. We believe that this is due to a lack of domain-specific pre-training. Our approach, in contrast, demonstrates the potential of pure transformer architectures for defect detection without architectural crutches like CNN layers.
3 Dataset
We use scanning acoustic microscopy (SAM) images and transient thermal analysis (TTA) measurements from a dataset of microscale solder joints of high-power LEDs [31], which is publicly available on Kaggle [30]. The dataset contains SAM and TTA data for 1.800 LEDs at 5 points of time, with nine LED types and five lead-free solder pastes. To simulate ageing, LED panels underwent thermal shock cycles (TSC) from $-40$°C to $125$°C, a standard quality control process to evaluate the reliability and durability of solder joints under extreme conditions. We use a train, validation and test split of 60%, 20% and 20% respectively. All reported metrics and all shown example outputs refer to the test dataset.
Table 1: Dataset. Repeated cooling and heating of LED solder joints simulates aging through thermomechanical fatigue. Over time, cracks gradually emerge, impeding the heat flow through the solder material. $\Delta B_{max}$ values indicate relative degradation and LEDs with $\Delta B_{max}>20\%$ are classified as defective ()(*)( ∗ ).
TSC 0 100 500 1.000 1.500
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$\Delta B_{max}$ $0\%$ $3\%$ $18\%$ $99\%$ ()(*)( ∗ ) $106\%$ ()(*)( ∗ )
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$\Delta B_{max}$ $0\%$ $2\%$ $30\%$ ()(*)( ∗ ) $62\%$ ()(*)( ∗ ) $75\%$ ()(*)( ∗ )
3.1 Scanning Acoustic Microscopy (SAM)
We use SAM images for pre-training and as model inputs in downstream tasks. SAM imaging offers detailed information on material interfaces, allowing the detection of critical defects such as voids, cracks, and delaminations within solder joints [23]. SAM images were taken after 0, 100, 500, 1.000, and 1.500 TSC, resulting in 5 images per LED and 9.000 images in total (see Table 1). The original SAM images contain 2 LEDs and are cropped so that each image shows an individual solder joint. All images are resized to a $64\times 64\times 1$ pixel format. For each LED image, bright boundary regions depict gaps between the darker solder pads. On the solder pads, small circles indicate voids in the solder material, from which cracks can emerge over time. Bright, non-circular structures inside the pads are cracks in the solder material (see Table 1).
3.2 Transient Thermal Analysis (TTA)
We use TTA data to create labels indicating how far LEDs depicted in SAM images are from failure. TTA is a non-destructive test method that quantifies the integrity of the thermal path from the LED junction to the heat sink through the solder joint [51]. Defects such as cracks impede the heat flow through the solder by reducing the size of the contact area. In TTA, the thermal impedance $Z_{th}(t)$ is measured, which describes the temperature difference at the LED junction over time compared to the change in power. Any degradation results in an increase in $Z_{th}(t)$. For a quantitative comparison of TTA measurements between different TSCs, the maximum value $B_{max}$ of the normalized logarithmic derivative $B(t^{\prime})$ of the thermal impedance $Z_{th}(t)$ with $t^{\prime}=ln(t)$ is used [52]. As $B_{max}$ increases in case of degradation, the relative increase in $B_{max}$ at cycle $t$ compared to the initial state $\Delta B_{max}(t)$ serves as the basis for image labelling:
$\Delta B_{max}(t)=\frac{B_{max}(t)}{B_{max}(0)}-1.$ (1)
While we use numerical labels to achieve a continuous representation of an LED’s current quality, an LED with $\Delta B_{max}>20\%$ is classified as defective [44].
4 Method
4.1 Preliminaries
4.1.1 Vision Transformer
Transformer processing employs multi-head self-attention (MSA) on sequential inputs [40]. For the sequentialization of 2D images, images are divided into smaller patches and are linearly projected into the embedding space. Let $H_{i},W_{i}$ be the height and width of the original image, and let $H_{p},W_{p}$ be the patch height and width, where usually image patches are quadratic, i.e. $H_{p}=W_{p}=:P$ (*). The image is then divided into
$N=\frac{H_{i}\cdot W_{i}}{H_{p}\cdot W_{p}}\overset{\textbf{$*$}}{=}\frac{H_{i}\cdot W_{i}}{P^{2}}$ (2)
patches. After flattening and applying a linear mapping $E\in\mathbb{R}^{P^{2}\times D}$ to project the patch into the embedding space $\mathbb{R}^{D}$, positional embeddings $E_{pos}\in\mathbb{R}^{N\times D}$ are added to the elements of the sequence to retain the spatial location of each patch embedding. Lastly, a learnable class token $x_{class}$ is prepended to the sequence to store global information during processing.Images are thus transformed into a sequence $X$ of $N+1$ embeddings, which serve as tokens for MSA:
$X=[x_{class},Ex_{1}+E_{pos,1},\dots,Ex_{N}+E_{pos,N}]\subseteq\mathbb{R}^{D}$ (3)
Depending on the ViT size, the sequence passes through several MSA blocks. Since we are working with relatively small $64\times 64\times 1$ shaped images, we use ViT sizes tiny, small, and base for determining the number of transformer blocks, the token length, and the number of heads (see Table  2). In the original ViT [13], the output class token is fed into a single linear layer for classification. We find that using a larger classification head increases performance and thus use a multi-layered dense network with a hidden dimension of $2048$.
Table 2: ViT Sizes. While all of the evaluated sizes use 12 transformer blocks, they differ in token length (width) and the number of heads used for MSA.
Model Layers Width Heads Params
ViT-Ti 12 192 3 5.7 M
ViT-S 12 384 6 22.1 M
ViT-B 12 768 12 86.7 M
4.1.2 Masked Autoencoder Pre-Training
For MAE pre-training, a ViT encoder combined with a lightweight transformer decoder is tasked with reconstructing missing image patches (see Figure 1). Depending on the masking share $S$, which is usually $0.75$, a random subset $M$ of $\lfloor N\cdot S\rfloor$ indexes is sampled from the set of indexes $I=\{1,\dots,N\}$ without replacement. $X$ is split into a set of masked patches $X_{M}$ and a set of visible patches $X_{I\setminus M}$. Note that $X_{M}$ and $X_{I\setminus M}$ are absolute complement with respect to $X_{I}$, i.e. $X_{m}\cup X_{I\setminus M}=X_{I}$ and $X_{m}\cap X_{I\setminus M}=\emptyset$. After masking, the sequence consists of $\lfloor N\cdot S\rfloor$ embeddings:
$X=[\tilde{x}_{1}E+E_{pos,1},\dots,\tilde{x}_{\lfloor N\cdot S\rfloor}E+E_{pos,\lfloor N\cdot S\rfloor}]\subseteq\mathbb{R}^{D}$ (4)
Inputs are then fed into the autoencoder network. The MAE encoder operates only on the visible patches, i.e. on a $1-S$ share of the total input patches. The decoder, in contrast, receives as input a full sequence of tokens including a) the embeddings of visible patches and b) mask tokens, which are shared and learnable embeddings serving as placeholders for masked patches, similar to BERT [12]. The model minimizes the pixel-wise MSE loss for masked patches with respect to its parameters $\theta$. The loss is computed for masked patches only, i.e. for $k\in M$:
$\min_{\theta}\text{ }\mathcal{L}_{mae}(X,M,\theta)=\sum_{k\in M}{\|y_{k}-x_{k}\|_{2}}^{2}$ (5)
Table 3: Hyperparameter tuning for patch size, mask ratio, and augmentations.
(a) Patch Sizes
Patch Size MSE $\downarrow$
$4\times 4$ 0.0186
8 $\times$ 8 0.0167
$16\times 16$ 0.0197
(b) Mask Ratios
Mask Ratio MSE $\downarrow$
70% 0.0175
75% 0.0167
80% 0.0188
(c) Augmentations
Augmentations MSE $\downarrow$
horizontal flip 0.0190
vertical flip 0.0189
both 0.0179
crop (random) 0.0185
crop (fixed) 0.0176
5 Experiments
5.1 Hyperparameter Tuning
Because of the different structure of SAM images and due to the smaller image size compared to standard datasets like ImageNet, we hypothesize that smaller patch sizes, different augmentations, and different mask ratios than those used for standard ViT could achieve optimal performance. To evaluate hyperparameter settings, we pre-train models with MAE for 200 epochs on our dataset, and fine-tune for defect detection for 100 epochs.
Results are shown in Table 3(c). Although we find that, consistent with the results for ImageNet, a mask ratio of $75\%$ works best, our hypothesis was confirmed for patch size and augmentations: For image patches, a size of $8\times 8$ worked best, while typically, a patch size of $16\times 16$ is used. Notably, the found optimal image to patch size ratio of $64:8=8$ also differs from the standard ratio of $224:16=14$ used for datasets like ImageNet [13]. We assume that this non-linear dependence between image size and optimal patch size is explained by the fact that once patches get too small, a single patch does not contain enough localized information for a meaningful computation of self-attention.
For augmentations, a combination of horizontal and vertical flipping and fixed-size cropping achieved the best results, while for ImageNet horizontal flipping and random-sized cropping are used [13]. In contrast to natural images, vertical flipping is a valid augmentation for microelectronics, where image orientation is not relevant. Also, we suspect that fixed-size cropping works best since random-sized crops of $64\times 64$ images suffer greater quality loss by resizing operations than the $224\times 224$ images in standard large datasets like ImageNet.
Figure 3: MAE Pre-Training.
5.2 MAE Self Pre-Training
We conduct MAE pre-training on the target SAM image dataset, utilizing the hyperparameters optimized through our previous studies. MAE with ViT encoders of sizes Ti, S, and B are pre-trained for 1.000 epochs on two L40S GPUs respectively. In Figure 3, we plot the MSE reconstruction loss for masked input patches $\mathcal{L}_{mae}$ for each model size. Notably, the S-sized encoder achieved the worst performance in reconstruction, although having more parameters than the Ti-sized encoder. However, the even larger B-sized encoder significantly outperformed all smaller sizes. We take this to indicate that a) for smaller images, even very small encoders like ViT-Ti can achieve good results and that b) there is a non-linear relationship between model capacity and performance. For smaller datasets like ours, there appears to be a critical threshold where increased model complexity transitions from potentially harmful (due to overfitting) to beneficial. Although ViT-B achieved the best results overall, ViT-Ti could be the model of choice in resource-critical applications.
Exemplary reconstructions from the test set using the best-performing ViT-B encoder are shown in Figure 4. We find that after training, the autoencoder produces high-quality reconstructions for all LED types, indicating that it learned meaningful representations of the dataset. In particular, the model is able to reconstruct masked inputs at varying amounts of TSC, accurately predicting the expansion of cracks even in masked regions (see Figure 4 (b)).
(a) 0 TSC
(b) 1500 TSC
Figure 4: MAE Reconstructions. For (a) TSC 0 and (b) TSC 1500, original images, masked inputs, and the model’s reconstructions are depicted.
5.3 Defect Detection
For inline automated optical inspection in microelectronics manufacturing, defect detection is a common practice. We perform defect detection given an input image using the $\Delta B_{max}$ values given by TTA as labels. While an LED with $\Delta B_{max}$ larger than $20\%$ is classified as faulty, we predict the scalar $\Delta B_{max}$ values directly to have a continuous measure for LED quality. That is, having the model predict $\Delta B_{max}$ not only indicates whether the given LED is functional, but also how far the LED is from failure. We compare ViT pre-trained with our MAE-based framework for 1.000 epochs against purely supervised ViT and ViT pre-trained with ImageNet. Also, we compare our models against large standard CNN backbones like ResNet [17], EfficientNet [36] or VGG [34] as well as against recent CNN architectures specifically designed for defect detection in microelectronics. In particular, we evaluate the model by Zippelius et al. [52] intended for SAM images of solder joints as well as the model by Zhang et al. [48] intended for X-ray images of solder joints. All models are trained for up to 200 epochs.
We find that ViT pre-trained with our self pre-training framework outperforms all other CNN and transformer-based models, achieving an MSE improvement of $5.30\%$ compared to the best-performing baseline (see Table 4). When it comes to other ViT models, our model outperforms purely supervised ViTs by $21.1\%$ and ViTs pre-trained on Imagenet by $10.2\%$. Notably, while the best results are achieved by the largest self pre-trained transformer model, ViT-Ti already outperforms considerably larger baselines by $3.20\%$ with only 5.7 million parameters. Thus, ViT self pre-trained with MAE not only substantially outperforms other ViT models and state-of-the-art CNN defect detection models, but is already able to do so using the smallest available architecture size.
Table 4: Defect Detection. All models are trained for 200 epochs on the SAM dataset. We report mean squared error (MSE) scores on the test set, where a lower score is better. Also, we report the percental differences to the best model.
Model Method Params MSE $\mathrm{\Delta}$ to Best
SAM-CNN [52] supervised 4.2 M 0.0336 10.5%
XRAY-CNN [48] supervised 8.7 M 0.0337 10.9%
ResNet50 [17] supervised 23.5 M 0.0339 11.5%
EfficientNet-B7 [36] supervised 63.8 M 0.0331 8.9%
VGG-11 [34] supervised 128.8 M 0.0320 5.3%
ViT-Ti [13] supervised 5.7 M 0.0380 25.0%
ViT-S [13] supervised 22.1 M 0.0368 21.1%
ViT-B [13] supervised 86.7 M 0.0380 25.0%
ViT-Ti [13] MAE (ImageNet) 5.7 M 0.0345 13.4%
ViT-S [13] MAE (ImageNet) 22.1 M 0.0339 11.5%
ViT-B [13] MAE (ImageNet) 86.7 M 0.0335 10.2%
ViT-Ti (ours) MAE (self) 5.7 M 0.0310 2.0%
ViT-S (ours) MAE (self) 22.1 M 0.0335 10.2%
ViT-B (ours) MAE (self) 86.7 M 0.0304 -
In contrast to our self pre-trained models however, we find that purely supervised ViT performs poorly compared to all other models due to the lack of pre-training and the small amount of target data, which is consistent with earlier findings [49]. What is more, we find that MAE pre-training using ImageNet increases performance compared to purely supervised ViT, but cannot outperform convolutional baselines. This suggests that for applications with even less data than in our case, fine-tuning foundational vision models or using convolutional defect detection architectures might remain a valid option.
We perform interpretability analysis with class activation maps using GradCAM [32]. In GradCAM, the relevance of image regions for the output is visualized by computing gradients of the output with respect to a given layer. We choose the first norm layer of the last transformer block for the gradient computation. Results for all ViT-B variants are shown in Figure 5. We find that:
  1. 1.
    Supervised ViT-B has scattered attention for both early and late TSC (see Figure 5), indicating that the transformer encoder does not learn meaningful representations in purely supervised training.
  2. 2.
    ViT-B pre-trained on ImageNet strongly focuses on the bright bounding regions between solder pads, indicating that the model attends to the LED type, even when clear defects such as cracks in the solder material are visible (see Figure 5 (b), row 3 and 4). We interpret this as the model learning a shortcut rather than the actual dependency between defects like cracks in the solder material and $\Delta B_{max}$.
  3. 3.
    ViT-B self pre-trained on SAM data solely attends to actual defects such as cracks for later TSC with stronger degradation (Figure 5 (b)) and attends to voids and the boundary regions between solder pads if no damages are visible (Figure 5 (a)). This indicates that the model prioritizes actual defects and considers the general error-proneness of different LED types when no defects are visible.
Thus, our findings indicate that MAE self pre-training yields superior representations for defect detection compared to supervised training and pre-training on natural image data. This confirms our hypothesis that domain-specific pre-training can account for the data requirements of transformers under the data sparsity in microelectronics.
(a) 0 - 500 TSC
(b) 1.000 - 1.500 TSC
Figure 5: Attention Visualization. Compared to supervised ViT-B and ViT-B pre-trained with MAE on ImageNet, our model self pre-trained with MAE focuses on the LED type and on voids in the solder material for earlier TSC (a), and on defects such as cracks for later TSC (b), indicating that MAE self pre-training yields superior and defect-specific feature representations.
6 Discussion and Conclusion
In this paper, we employ vision transformers self pre-trained using masked-autoencoders for defect detection in microelectronics. Our methodology leverages the strong predictive capabilities of transformers while adapting to our specific target domain despite limited labeled data. For defect detection on a small dataset of microscale solder joints of high-power LEDs, our approach demonstrates superior performance compared to various state-of-the-art CNN-based architectures and other transformer baselines.
The demonstrated effectiveness of our approach has broad implications for quality control in Industry 4.0 settings, where maximizing production yield while minimizing the cost and time required for inspection remains a central challenge. By enabling more accurate detection of developing solder joint defects in early stages, our approach can reduce warranty costs, improve product reliability, and increase customer satisfaction. As manufacturing processes continue to demand higher precision and reliability while dealing with ever-increasing miniaturization, approaches like ours that can leverage limited domain data effectively will become increasingly valuable. The combination of high accuracy, interpretability, and resource efficiency positions our MAE self pre-training framework as a promising path forward for next-generation quality inspection systems in microelectronics, with particular potential for integration into capable smart sensors that enable real-time defect detection directly on the production line.
{credits}
6.0.1 Acknowledgements
The presented work is part of the project MaWis-KI, which is supported by the Bavarian Ministry of Economic Affairs, Regional Development and Energy in the context of the Bavarian funding program for research and development "Information and Communication Technology".
[…]
The authors have no competing interests to declare that are relevant to the content of this article.